Solve by I.F method


-2xydy+(x2+y2)dx = 0


1
Expert's answer
2021-11-24T16:14:58-0500

"-2xydy+(x^2+y^2)dx = 0"


"2yy'-y^2\/x=x"

"v=y^2"

"v'-v\/x=x"


integrating foctor:

"\mu=1/x"


"v'\/x-v\/x^2=1"


"v'/x+\frac{d}{dx}(1/x)v=1"


"\frac{d}{dx}(v/x)=1"


"v\/x=x+c"


"y=\pm \sqrt{x(x+c)}"